OpenStudy (anonymous):

A rectangular prism has a length of 12 cm and a width of 4 cm. If the volume of the prism is 240 cm3, what is the height? 3 cm 4 cm 5 cm 6 cm

OpenStudy (anonymous):

vol = 12 * 4 * h = 240 h = 240 / 48 = 5 cm